OpenStudy (anonymous):

passing through (-15, -13) and parallel to the line whose equation is y=-8x-4 put in point slope form and slope intercept form

OpenStudy (dmndlife24):

One sec

OpenStudy (dmndlife24):

y=-8x-133

OpenStudy (anonymous):

thank you! how do i find point slope form?

OpenStudy (dmndlife24):

-15 is x1 -13 is y1

OpenStudy (anonymous):

oh i see it

OpenStudy (dmndlife24):

point slope is y-y1=m(x-x1)

OpenStudy (dmndlife24):

m = -8

OpenStudy (anonymous):

i wrote =-8(x-(-15) for the second part and it says its wrong?

OpenStudy (dmndlife24):

Did you simplify?

OpenStudy (anonymous):

yea said its wrong too

OpenStudy (anonymous):

whoops I was putting a 'x' at the end of the 8